What is the correct procedure to solve The mixture calls for 75% pure water and 25% concentrated vitamin drink. How many gallons of water should you add to 16 gallons of concentrated vitamin drink?
Found 2 solutions by ewatrrr, Theo:
Answer by ewatrrr(24785)

 
Hi,
The mixture calls for 75% pure water and 25% concentrated vitamin drink.
mixture of water to concentrate is 3:1
48gal of water would need to be added to 16gal of concentrate

Answer by Theo(13342)

one way to solve this is as follows:
75% of the mixture is water.
25% of the mixture is concentrate.
since there are 16 gallons of concentrate, this means that 16 is equal to 25% of the mixture.
if that's true, then the mixture must be equal to 4 * 16 = 64.
since water is 75% of the mixture, this means that water must be equal to 64 - 16 = 48.
you have 48 gallons of water and 16 gallons of concentrate to equal a total of 64 gallons of mixture.
48/64 = .75
16/64 = .25

algebraically this looks like the following:
let w = number of gallons of water.
let c = number of gallons of concentrate
let m = number of gallons of the mixture.
you are given that:
c = 16
you are also given that:
c = .25*m
this means that:
16 = .25*m
divide both sides of this equation by .25 to get:
16/.25 = m
simplify to get:
m = 64
you are given that:
w = .75*m
replace m with 64 to get:
w = .75*64
simplify to get:
w = 48

another way to solve is:
let w = number of gallons of water.
let c = number of gallons of concentrate
let m = number of gallons of the mixture.
you are given:
c = .25*m
w = .75*m
from the first equation, solve for m to get:
m = c/.25
replace m in the second equation with c/.25 to get:
w = .75*c/.25
simplify to get:
w = 3*c
since c = 16, this means that m = 48.